OpenStudy (anonymous):

What is 1/2 divided by 3/10

OpenStudy (anonymous):

5/3

OpenStudy (angela210793):

1/2*10/3=10/6=5/3

OpenStudy (angela210793):

yea LOL :P

OpenStudy (anonymous):

did u understand jjsonfire ?